# PAID: A Framework of Product-Centric Advertising Image Design

### v1

@@ 5.4. Ablation Studies
Effect of RKBR and CCLP in layout generation. We validate the effectiveness of the proposed input and output format for layout generation, including RKBR and CCLP. As shown in Table 6, without RKBR, the model may predict a product size which does not match the original aspect ratio about 5% of the time. Note that we regard a difference of less than 1.5% as correctness. Adding CCLP improves the Uti and Occ values on No Occ Set, indicating that CCLP helps the model to distinguish occlusion-allowing sets and others by providing explicit class conditions.
@@
Effect of DLC and LAT in layout-to-image model. We investigate the effect of deep layer control (DLC) and Lora adaptation training (LAT), with results shown in Table 7. Adding layout control in deep UNet layers performs almost as well as controlling all layers. It reduces parameters and inference costs, which benefits the application. With LoRA adaptation, the model achieves lower FID and higher mIoU, enhancing image quality and spatial control. Parameters and inference costs can be further reduced.
@@
6. Conclusion
In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create ad images only based on product foreground images, marketing taglines, and target sizes. PAID consists of four serial stages: prompt generation, layout generation, background image generation, and graphics rendering. First, it generates adaptive prompts according to the semantic and visual information of the product, for the benefit of background image generation. Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout. Next, a layout-controlled inpainting model is utilized for background image generation. A graphics rendering module is then applied to get the final images. We train separate experts to conduct these sub-tasks. Two high-aesthetic and diverse product-centric datasets are created for the convenience of training. Qualitative and quantitative comparisons prove that the proposed PAID framework generates more visually pleasing ad images, compared with previous methods.
Acknowledgements

### v2

@@ C.1. Effect of JPGNL in Layout Generation
C.2. Effect of RKBR and CCLP in Layout Generation
We validate the effectiveness of the proposed input and output format for layout generation, including RKBR and CCLP. As shown in Table 6, without RKBR, the model may predict a product size which does not match the original aspect ratio about 5% of the time. Note that we regard a difference of less than 1.5% as correctness. Adding CCLP improves the Uti and Occ values on No Occ Set, indicating that CCLP helps the model to distinguish occlusion-allowing sets and others by providing explicit class conditions.
@@ C.2. Effect of RKBR and CCLP in Layout Generation
C.3. Effect of DLC and LAT in Layout-to-Image Model
We investigate the effect of deep layer control (DLC) and Lora adaptation training (LAT), with results shown in Table 7. Adding layout control in deep UNet layers performs almost as well as controlling all layers. It reduces parameters and inference costs, which benefits the application. With LoRA adaptation, the model achieves lower FID and higher mIoU, enhancing image quality and spatial control. Parameters and inference costs can be further reduced.